OpenStudy (anonymous):

integral of ln(x)^3 dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\ln(x)^3 dx\]

OpenStudy (anonymous):

Sorry, \(ln(x^3)\) or \((ln(x))^3\) ??

OpenStudy (anonymous):

\[\ln(x)^3\]

OpenStudy (anonymous):

i know how to do \[(lnx)^3\]

OpenStudy (anonymous):

Use power rule then: \[\ln(a^b) = b \cdot \ln(a)\]

OpenStudy (abb0t):

@Somy tells me that only those who cant land jobs go into higher education after bachelors...

OpenStudy (anonymous):

@abb0t then what can we do in this??

OpenStudy (abb0t):

I am in islamabad

OpenStudy (anonymous):

Good for you.. :)

OpenStudy (abb0t):

Do you think @Somy is right about this?

OpenStudy (anonymous):

\[\int\limits \ln(x)^3 dx = 3 \int\limits \ln(x).dx\]

OpenStudy (anonymous):

No, upto some extent, he or she may be right..

OpenStudy (anonymous):

@yoonjae can you proceed further??

OpenStudy (anonymous):

how do u integrate ln x

OpenStudy (anonymous):

With the pen in your hand on your notebook.. :P

OpenStudy (anonymous):

use integration by parts ?

OpenStudy (anonymous):

Yep.. :) Good..

OpenStudy (anonymous):

are u sure about power rule?

OpenStudy (anonymous):

You can write it as: \[3\int\limits x^{0} \cdot \ln(x) \cdot dx\]

OpenStudy (anonymous):

I think I am...

OpenStudy (anonymous):

If this is \(ln(x^3)\) then I am sure..

OpenStudy (anonymous):

so i got 3xlnx -3x + C

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

Good, you must get that only.. :)

OpenStudy (anonymous):

thank you sir

OpenStudy (anonymous):

"Friend"..

OpenStudy (anonymous):

okay bro =))

OpenStudy (anonymous):

Okay brother.. :)

OpenStudy (abb0t):

uck.

OpenStudy (somy):

lol did he delete his answer? (i mean dan (from my acc))